OpenStudy (anonymous):

Find the doubling time of a quantity that is increasing 8% per year. Round your answer to two decimal places. The doubling time is roughly _______ years.

OpenStudy (anonymous):

I used: P(t) = Poe^rt ---> P(t)=2Po ---> 2Po=Poe^rt 2=e^rt ln(2)=rt t=ln(2)/0.08 = 8.66 years What did I do wrong?

OpenStudy (anonymous):

Figured it out: t = log 2/log 1.08 = 9.006468 years